### v1

# Off-Equatorial Orbits around Magnetically Charged Black Holes

Abstract
We present a complete characterization of stable, off-equatorial circular orbits around magnetically charged black holes (MBHs). For a static, spherically symmetric MBH, we derive an exact analytic expression for the orbital latitude $\theta$ as a function of radius $r$. We establish a direct connection between these orbits and the spacetime fundamental structure, and demonstrate their stability against synchrotron radiation. We show that charged particles such as electrons and protons can exhibit $\mathcal{O}(1)$ latitude deviations at the ISCO radius and remain stable under synchrotron emission even for extremely small values of the black hole magnetic charge. We then extend the analysis to rotating MBHs, numerically computing the prograde and retrograde orbital branches and demonstrating how frame-dragging modifies their structure and stability regions. We show that these off-equatorial orbits are a unique feature of the magnetic charge, being forbidden in the analogous electrically charged Kerr-Newman spacetime. Our results suggest that environments surrounding magnetically charged black holes can exhibit distinctive phenomenological signatures, with potential implications for black hole imaging and polarimetric observations.
I Introduction
The coupling of magnetic charge to gravity provides a rich theoretical laboratory for probing classical and semiclassical black-hole physics [MBH1, MBH2, Bai:2020spd, Bai:2020ezy, Gervalle:2024yxj]. The concept of magnetic monopole arises naturally in many extensions of the Standard Model, particularly in grand-unified theories (GUTs), where they are predicted to form as topological defects during symmetry-breaking phase transitions in the early universe [Dirac1931, tHooft1974, Polyakov1974]. Their gravitational counterparts, magnetically charged black holes (MBHs), serve as an idealized framework for studying how non-trivial gauge fields alter spacetime geometry and the dynamics within it, including the causal structure, particle motion, and radiative processes [Gibbons1975, Lee1992, Tursunov2018, Rayimbaev2021]. A key feature of these spacetimes is the non-zero divergence of the magnetic field, which prevents the electromagnetic vector potential from being defined globally. Nevertheless, a local potential can be defined in regions free of the monopole source, allowing for a consistent analysis of charged particle dynamics [Rohrlich1966, Preskill1984]. Furthermore, MBHs are connected to their electrically charged cousins—the Reissner–Nordström and Kerr–Newman black holes—via electromagnetic duality, a fundamental symmetry of Maxwell’s equations where the roles of electric and magnetic fields can be interchanged [duality1, duality2]. Studying the phenomenology of MBHs is therefore of intrinsic theoretical interest and may also inform searches for hypothetical observational signatures of such compact objects [Liu:2020vsy, Liu:2020bag, Chen:2022qvg] or distinguish them from exotic mimickers [Bah:2025vbr].
The analysis of particle and light trajectories around black holes is a cornerstone of relativistic astrophysics [orbit1, orbit2, orbit3]. The paths of light are described by null geodesics, which determine the properties of gravitational lensing and define the location of the photon sphere—a region composed of unstable circular light orbits that is fundamental to the appearance of a black hole’s shadow [Synge1966, Bardeen1973, Chael2021]. The paths of massive objects are described by timelike geodesics, which govern the dynamics of accretion disks and define the innermost stable circular orbit (ISCO). The ISCO represents the minimum radius at which a particle can stably orbit a black hole before beginning an inevitable plunge, making it a critical parameter in models of accretion efficiency and compact-object inspirals [Bardeen1972, CompactInspiral, ISCO1, ISCO2]. The dynamics become significantly richer when the test particles are electrically charged. The Lorentz force, acting in concert with gravity, can produce qualitatively novel phenomena, such as strong latitudinal deflection and the emergence of stable, off-equatorial orbits not possible for neutral particles [Carter1968, Pugliese2011, Hackmann2013, Russo:2020lah, Dyson:2023ujk, Pereniguez:2024fkn]. A comprehensive analytic characterization of circular bound orbits around static and rotating MBHs, together with an analysis of their stability under electromagnetic emission and a connection between their properties and the fundamental structure of the spacetime, is still lacking in the literature. The present work aims to fill this gap.
In this paper, we analyze the orbital structure of neutral and charged test particles in two key backgrounds, adopting natural units where $c=G=1/4\pi\epsilon_{0}=1$. The first is the static, spherically symmetric MBH, described by the metric function
$f(r)=1-\frac{2M}{r}+\frac{P^{2}}{r^{2}},$ (1)
[…]
II Orbital Structure in Static MBH
We begin with the static, spherically symmetric MBH, whose geometry is described by the line element in Schwarzschild-like coordinates $(t,r,\theta,\phi)$:
$ds^{2}=-f(r)\,dt^{2}+f(r)^{-1}dr^{2}+r^{2}d\Omega^{2},$ (2)
where the metric function $f(r)$ is given by Eq. (1) and $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$ is the line element on the unit 2-sphere. This metric is the magnetic analogue of the Reissner–Nordström spacetime and constitutes an exact solution to the Einstein–Maxwell equations for a purely radial magnetic field $\mathbf{B}=(P/r^{2})\,\hat{\mathbf{r}}$ [Reissner1916]. The cosmic censorship hypothesis imposes the bound $|P|\leq M$ [Penrose1965, Krolak1986], for which the spacetime possesses an event horizon ($r_{+}$) and an inner Cauchy horizon ($r_{-}$) at
$r_{\pm}=M\pm\sqrt{M^{2}-P^{2}}.$ (3)
In the extremal limit, $|P|\to M$, these horizons coincide at $r_{\pm}=M$.
Due to the magnetic charge, the electromagnetic four-potential $A_{\mu}$ cannot be globally defined. It can, however, be expressed locally on gauge patches. For the northern hemisphere ($0\leq\theta<\pi$), we adopt the standard form [Rohrlich1966, Preskill1984]:
$A^{(N)}_{\phi}=P(1-\cos\theta),\qquad A^{(N)}_{t}=A^{(N)}_{r}=A^{(N)}_{\theta}=0.$ (4)
The potential on the southern hemisphere, $A^{(S)}$, is related by a gauge transformation, $A^{(N)}-A^{(S)}=2P\,d\phi$. Henceforth, we work in the northern hemisphere and omit the superscript $(N)$. The only non-vanishing components of the covariant electromagnetic tensor $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ are $F_{\theta\phi}=-F_{\phi\theta}=P\sin\theta$.
The motion of a test particle with rest mass $m$ and electric charge $q$ is governed by the the geodesic equation, for a neutral particle ($q=0$) it gives:
$\frac{d^{2}x^{\mu}}{d\tau^{2}}+\Gamma^{\mu}_{\alpha\beta}u^{\alpha}u^{\beta}=0.$ (5)
For a charged particle, the equation of motion includes the Lorentz force:
$\frac{d^{2}x^{\mu}}{d\tau^{2}}+\Gamma^{\mu}_{\alpha\beta}u^{\alpha}u^{\beta}=\frac{q}{m}F^{\mu\nu}u_{\nu},$ (6)
where $\Gamma^{\mu}_{\alpha\beta}$ are the Christoffel symbols.
For neutral particles, the spacetime symmetries guarantee the conservation of energy $E$ and angular momentum $J$. For photons ($m=0,q=0$), which satisfy the null condition $g_{\mu\nu}u^{\mu}u^{\nu}=0$, circular orbits define the photon sphere. Its radius, $r_{\text{ps}}$, is found by extremizing the effective potential, yielding
$r_{\text{ps}}=\frac{3M\pm\sqrt{9M^{2}-8P^{2}}}{2}.$ (7)
where $r_{\text{ps}}^{+}$ and $r_{\text{ps}}^{-}$ correspond to the outer and inner photon spheres, respectively. The outer sphere represents the physically relevant photon sphere, and its radius decreases monotonically from the Schwarzschild value of $3M$ to $2M$ as $|P|$ increases from 0 to $M$.
For massive neutral particles ($m\neq 0,q=0$), the innermost stable circular orbit (ISCO) marks the boundary between stable and plunging trajectories. The ISCO condition in the magnetically charged spacetime parallels that of the Reissner–Nordström case. The effective potential is given by
$V_{\text{eff}}=f(r)\left(1+\frac{j^{2}}{r^{2}}\right),$ (8)
where $j$ is the angular momentum per unit mass of the particle. The ISCO is determined by the standard marginal stability conditions,
$\frac{dV_{\text{eff}}}{dr}=\frac{d^{2}V_{\text{eff}}}{dr^{2}}=0.$ (9)
which yields the cubic equation
$Mr^{3}-6M^{2}r^{2}+9MP^{2}r-4P^{4}=0.$ (10)
The real root of Eq. (10) gives the ISCO radius $r_{\text{ISCO}}$, which decreases monotonically from the Schwarzschild value $r_{\text{ISCO}}=6M$ to $4M$ in the extremal limit.
The presence of electric charge of the test particle introduces a Lorentz force that alters the orbital plane, displacing the trajectory from the equator. Stable circular orbits still exist for certain polar angles, as shown in Figure 1.
Figure 1: Circular orbit of a charged particle around an MBH. The magnetic force shifts the particle to a nontrivial polar angle.
For a test particle moving on a circular orbit, its 4-velocity is given by
$u^{\mu}=(u^{t},0,0,u^{\varphi}),$ (11)
with the constraint
$u^{\mu}u_{\mu}=-f(r)(u^{t})^{2}+\frac{\dot{r}^{2}}{f(r)}-r^{2}(u^{\varphi})^{2}=-1.\\$ (12)
The $t$ and $\varphi$ components of Eq. (6) vanish identically as
$\begin{split}&\Gamma^{t}_{tt}=\Gamma^{t}_{t\varphi}=\Gamma^{t}_{\varphi\varphi}=0,\\ &\Gamma^{\varphi}_{tt}=\Gamma^{\varphi}_{t\varphi}=\Gamma^{\varphi}_{\varphi\varphi}=0,\\ &F^{tt}=F^{t\varphi}=F^{\varphi t}=F^{\varphi\varphi}=0.\end{split}$ (13)
The $r$ and $\theta$ components then determine the polar angle and angular velocity for a given radius. We begin with the $r$-component, which has no Lorentz force term. Using the relevant Christoffel symbols
$\begin{split}&\Gamma^{r}_{t\varphi}=0,\\ &\Gamma^{r}_{tt}=(1-\frac{2M}{r}+\frac{P^{2}}{r^{2}})(\frac{M}{r^{2}}-\frac{P^{2}}{r^{3}}),\\ &\Gamma^{r}_{\varphi\varphi}=-r\sin^{2}\theta(1-\frac{2M}{r}+\frac{P^{2}}{r^{2}}),\end{split}$ (14)
we obtain the angular velocity
$\frac{d\varphi}{d\tau}=\omega_{\pm}=\pm\frac{1}{r\sin\theta}\sqrt{\frac{\frac{M}{r}-\frac{P^{2}}{r^{2}}}{1-\frac{3M}{r}+\frac{2P^{2}}{r^{2}}}}.$ (15)
where $\omega_{+}$ and $\omega_{-}$ correspond to prograde (meaning counterclockwise as viewed from above the northern hemisphere) and retrograde motion (opposite), respectively.
For the $\varphi$ equation, the electromagnetic contribution yields
$F^{\mu\nu}u_{\nu}=g^{\mu\alpha}g^{\nu\beta}F_{\alpha\beta}u_{\nu}=\frac{P}{r^{2}}\omega\sin\theta,$
while the geodesic term is
$\Gamma^{\varphi}_{\alpha\beta}\frac{dx^{\alpha}}{d\tau}\frac{dx^{\beta}}{d\tau}=\Gamma^{\varphi}_{\varphi\varphi}u^{\varphi}u^{\varphi}=-\omega\sin\theta\cos\theta.$
Combining these, the only possible circular orbit for the charged particle lies at
$\tan(\theta_{\pm})=\mp\frac{mr}{qP}\sqrt{\frac{\frac{M}{r}-\frac{P^{2}}{r^{2}}}{1-\frac{3M}{r}+\frac{2P^{2}}{r^{2}}}}.$ (16)
These orbits exist only when the quantity under the square root is positive, which requires the radius to be larger than the photon-ring radius, as will be shown in Section III. For a charged particle located at $r_{\text{ISCO}}$, Eq. (16) is accurately approximated by the expression
$\tan(\theta_{\pm})=\mp\,\,2\sqrt{3}\,\,\bigg(\frac{q}{m}\bigg)^{\!\!-1}\!\!\left(\frac{P}{M}\right)^{\!\!-1}\,.$ (17)
This expression indicates that a realistic particle, such as an electron with a charge-to-mass ratio of $\sim 10^{22}$, can still have $\tan(\theta)\sim\mathcal{O}(1)$ for values of $P/M$ as small as $10^{-22}$, highlighting the strong influence that even a tiny magnetic charge can exert on light charged particle. This is somehow reminiscent of the findings in Ref. [Pereniguez:2024fkn], where the authors show that even a single magnetic monopole in the black hole can significantly affect the superradiant instability timescale.
Figure 2 shows $\theta_{+}(r)$ for a q/m=1q/m=-1 particle and various magnetic charge-to-mass ratios of the MBH. Here $\theta_{+}<\pi/2$ because the magnetic force on a prograde, negatively charged particle is directed above the equator, decreasing its polar angle. From Eq. (16), such orbits exist only for $r$ larger than a certain threshold. The angle approaches $\pi/2$ in both the small- and large-$r$ limits, corresponding to equatorial motion. There is always a single peak in $\theta(r)$, whose location will be shown in Section III to coincide with the ISCO radius, while the smallest orbital radius matches the photon-sphere radius. The stability of these orbits against radiative decay will also be considered in Section III.
Figure 2: $\theta_{+}(r)$ of a q/m=1q/m=-1 particle, for multiple $P/M$ MBHs. $\theta_{+}<\pi/2$ since the magnetic force on a prograde, negatively charged particle points above the equator. The y-axis is inverted so that north pole direction ($\theta=0$) appears at the top of the figure, matching the intuitive orientation.
Although $\theta\to\pi/2$ as $r\to\infty$, the particle does not necessarily approach the equatorial plane. The vertical distance from the plane, $y=r|\cos\theta|$, can be found from Eq. (16):
$y=r|\cos\theta|=\frac{r}{\sqrt{\tan^{2}\theta+1}}.$ (18)
In the large-$r$ limit, the asymptotic behavior of $y$ is
$y\sim\frac{r}{\sqrt{\frac{m^{2}r^{2}}{q^{2}P^{2}}\frac{M}{r}+1}}\sim\sqrt{r}\to\infty.$ (19)
Thus, the orbit remains at an asymptotically diverging vertical distance from the equatorial plane, despite that $\theta\to\pi/2$.
Figure 3 shows the upper half of a cross section of a MBH with $P/M=0.6$, together with representative prograde circular orbits for test particles with q/m=1q/m=-1 and q/m=10q/m=-10. The curves in the $(r,\theta)$ plane indicate a continuous series of such orbits. The black disk marks the horizon, while the grey and blue dashed circles denote the photon sphere and the ISCO radius, respectively. This is a striking consequence of the long-range magnetic field.
Figure 3: Cross-sectional view of circular orbits around a magnetically charged black hole with $P/M=0.6$. The black disk marks the event horizon, while the grey and blue dashed circles denote the photon sphere and the ISCO of neutral particles, respectively. The orange and purple curves indicate a series of orbits of prograde particles with q/m=1q/m=-1 and q/m=10q/m=-10.
III Analysis of Orbits in the Static Spacetime
We now perform a detailed analysis of the off-equatorial circular orbits derived in the previous section, connecting their properties to the fundamental features of the spacetime and assessing their stability.
The existence of the charged-particle orbits described by Eq. (16) depends on the reality of the square-root term. For radii outside the event horizon ($r>r_{+}$), the numerator ($M/r-P^{2}/r^{2}$) is positive. The denominator, however, vanishes at
$r=\frac{3M+\sqrt{9M^{2}-8P^{2}}}{2},$
which is precisely the photon-sphere radius given in Eq. (7). For a stable orbit to exist, the denominator must be positive, which requires $r>r_{\text{ps}}$. Thus, the photon sphere marks the innermost boundary for these off-equatorial circular orbits. As an orbit approaches this boundary from above ($r\to r_{\text{ps}}^{+}$), Eq. (16) shows that $\tan(\theta)\to\infty$, implying $\theta\to\pi/2$. This has a clear physical interpretation: at the photon sphere, the particle’s velocity approaches the speed of light, and its trajectory becomes equatorial, mimicking a null geodesic.
Next, we turn to the innermost stable circular orbit (ISCO) of a charged particle. Unlike the neutral case, the Lorentz force alters the direction of the tangential velocity, so the orbit cannot remain confined to the equatorial plane. Thus, the 4-velocity must be written as
$u^{\mu}=(\dot{t},\dot{r},\dot{\theta},\dot{\varphi}),$
and the normalization condition becomes
$u^{\mu}u_{\mu}=\frac{\dot{r}^{2}}{f(r)}+r^{2}\dot{\theta}^{2}+r^{2}\sin^{2}(\theta)\dot{\varphi}^{2}-f(r)\dot{t}^{2}=-1.\\$ (20)
The corresponding particle Lagrangian is
$\mathcal{L}=-\frac{m}{2}u^{\mu}u_{\mu}+qu^{\mu}A_{\mu}=-\frac{m}{2}u^{\mu}u_{\mu}+qA_{\varphi}\dot{\varphi}.$ (21)
Time independence of the Lagrangian guarantees conservation of energy:
$E_{\text{particle}}=-\frac{\partial\mathcal{L}}{\partial u^{t}}=-mu_{t}=mf(r)\dot{t}.\\$ (22)
Working per unit rest mass, we define
$\varepsilon=\frac{E_{\text{particle}}}{m}=mu_{t}=f(r)\dot{t},\\$ (23)
so that the temporal term in Eq. (20) becomes $\varepsilon^{2}/f(r)$.
The angular terms require more care, as neither $u^{\theta}$ nor $u^{\varphi}$ is separately conserved. However, the combination representing the total angular momentum magnitude per unit mass, is found to be conserved. To prove it, we define
$j=u^{\theta}u_{\theta}+u^{\varphi}u_{\varphi}=r^{2}[\dot{\theta}^{2}+\sin^{2}(\theta)\dot{\varphi}^{2}].$ (24)
The total derivative of $j$ is:
$\frac{dj}{d\tau}=\dot{g}_{\theta\theta}(u^{\theta})^{2}+2g_{\theta\theta}u^{\theta}\dot{u}^{\theta}+\dot{g}_{\varphi\varphi}(u^{\varphi})^{2}+2g_{\varphi\varphi}u^{\varphi}\dot{u}^{\varphi}.$ (25)
where we have
$\begin{split}&\dot{g}_{\theta\theta}=\frac{d(r^{2})}{d\tau}=2r\dot{r},\\ &\dot{g}_{\varphi\varphi}=\frac{d(r^{2}\sin^{2}(\theta))}{d\tau}=2r\dot{r}\sin^{2}(\theta)+r^{2}\sin(2\theta)\dot{\theta},\\ &\dot{u^{\theta}}=\frac{d^{2}u^{\theta}}{d\tau^{2}}=-\Gamma^{\theta}_{\mu\nu}u^{\mu}u^{\nu}+\frac{q}{m}F^{\theta\varphi}u_{\varphi},\\ &\dot{u^{\varphi}}=\frac{d^{2}u^{\varphi}}{d\tau^{2}}=-\Gamma^{\varphi}_{\mu\nu}u^{\mu}u^{\nu}+\frac{q}{m}F^{\varphi\theta}u_{\theta}.\end{split}$ (26)
Expanding Eq. (25) by Eq. (26), we get
$\begin{split}\frac{dj}{d\tau}=&2r\dot{r}\dot{\theta}^{2}+2r^{2}\dot{\theta}\left(-\Gamma^{\theta}_{\mu\nu}u^{\mu}u^{\nu}+\frac{q}{m}F^{\theta\varphi}r^{2}\sin^{2}(\theta)\dot{\varphi}\right)\\ &+dot{\varphi}^{2}\left[2r\dot{r}\sin^{2}(\theta)+2r^{2}\sin(\theta)\cos(\theta)\dot{\theta}\right]\\ &+2r^{2}\sin^{2}(\theta)\dot{\varphi}\left(-\Gamma^{\varphi}_{\mu\nu}u^{\mu}u^{\nu}+\frac{q}{m}F^{\varphi\theta}r^{2}\dot{\theta}\right).\end{split}$ (27)
Note that the 2 terms involving the electromagnetic tensor cancel out since $F^{\varphi\theta}=-F^{\theta\varphi}$. The non-vanishing Christoffel symbols are
$\begin{split}&\Gamma^{\varphi}_{\varphi r}=\Gamma^{\varphi}_{r\varphi}=1/r,\quad\Gamma^{\varphi}_{\theta\varphi}=\Gamma^{\varphi}_{\varphi\theta}=cot(\theta),\\ &\Gamma^{\theta}_{\varphi\varphi}=-\sin(\theta)\cos(\theta),\quad\Gamma^{\theta}_{\theta r}=\Gamma^{\theta}_{r\theta}=1/r.\end{split}$ (28)
Plugging them into Eq. (27), we finally obtain
$\begin{split}\frac{dj}{d\tau}=&r^{2}\dot{\theta}\left(\sin(\theta)\cos(\theta)\dot{\varphi}^{2}-\frac{2}{r}\dot{\theta}\dot{\varphi}\right)+2r\dot{r}\sin^{2}(\theta)\dot{\varphi}^{2}\\ &+2r\dot{r}\dot{\theta}^{2}+r^{2}\sin(\theta)\cos(\theta)\dot{\theta}\dot{\varphi}^{2}-2r\dot{r}\sin^{2}(\theta)\dot{\varphi}^{2}\\ &-2r^{2}\sin(\theta)\cos(\theta)\dot{\theta}\dot{\varphi}^{2}=0.\end{split}$ (29)
which proves the conservation of $j$.
Consequently, the radial equation reduces to
$\dot{r}^{2}=\varepsilon^{2}-f(r)(1+\frac{j^{2}}{r^{2}}),$
with effective potential identical to Eq. (8):
$V_{\text{eff}}=f(r)(1+\frac{j^{2}}{r^{2}}).$ (30)
The ISCO radius is then determined by the usual condition Eq. (10), identical to the neutral case. Thus, the particle charge does not shift the ISCO radius itself; rather, it controls the latitude of the orbit.
To solidify this connection, we can show that the extremum of the orbital latitude $\theta(r)$ occurs at the ISCO. The condition for an extremum is $d\theta/dr=0$, which is equivalent to $d(\tan\theta)/dr=0$. Applying this to Eq. (16) gives
$\frac{d}{dr}\left(r\sqrt{\frac{\frac{M}{r}-\frac{P^{2}}{r^{2}}}{1-\frac{3M}{r}+\frac{2P^{2}}{r^{2}}}}\right)=0.$ (31)
Evaluating the derivative yields the condition
$\frac{Mr^{3}-6M^{2}r^{2}+9MP^{2}r-4P^{4}}{r^{3}}=0.$ (32)
The numerator is exactly the polynomial from the ISCO condition in Eq. (10). Therefore, the maximum latitudinal deviation for a charged particle’s orbit occurs precisely at the ISCO radius of a neutral particle.
III.1 Synchrotron radiation of charged particles
Finally, we analyze the effect of synchrotron radiation on circular orbits, assessing the conditions under which these orbits can persist for sufficiently long lifetimes. In curved spacetime, the equation of motion of a charged particle incorporating the radiation-reaction effects reads [Bracken2024]
$m\ddot{z}^{\alpha}$ $=qF^{\alpha}_{\beta}\dot{z}^{\beta}+\frac{2}{3}q^{2}(\dddot{z}^{\alpha}-\ddot{z}^{2}\dot{z}^{\alpha})-\frac{1}{3}q^{2}(R^{\alpha}_{\beta}\dot{z}^{\beta}$
$+\dot{z}^{\alpha}R_{\beta\gamma}\dot{z}^{\beta}\dot{z}^{\gamma})+q^{2}\dot{z}^{\beta}\int^{\tau}_{-\infty}\zeta^{\alpha}_{\beta\gamma}\ddot{z}^{\gamma}(\tau^{\prime})d\tau^{\prime}$
$\equiv qF^{\alpha}_{\beta}\dot{z}^{\beta}+f^{\alpha}_{\text{self}}\,$, (33)
where $z(\tau)$ is the particle trajectory, and overdots denote covariant derivatives with respect to proper time $\tau$ (giving the 4-velocity, 4-acceleration, etc.). The first term of the right-hand side is the usual Lorentz force, while $f^{\alpha}_{\text{self}}$ denotes the self-force for a generic curved background.
The existence of a timelike Killing vector $k^{\mu}=(1,0,0,0)$ allows us to define the energy of the particle as $E=-m\,k_{\mu}\dot{z}^{\mu}$. The emitted power is then found by computing
$\dot{E}=-k_{\mu}(m\ddot{z}^{\mu})\,$, (34)
where we used the fact that $k^{\mu}$ satisfies the Killing equation $\nabla_{\alpha}k_{\mu}+\nabla_{\mu}k_{\alpha}=0$.
Using Eq. (33), the total power of radiation is finally expressed as
$P_{\text{tot}}=-\dot{E}=k_{\mu}f^{\mu}_{\text{self}}=(f_{\text{self}})_{0}\,.$ (35)
As expected, the Lorentz force yields no contribution to the emitted power.
Eq. (33) involves the evaluation of the following non-local term,
$q^{2}\dot{z}^{\beta}\int^{\tau}_{-\infty}\zeta_{0\,\beta\gamma}\ddot{z}^{\gamma}(\tau^{\prime})d\tau^{\prime}\,$, (36)
arising from backscattering of radiation off the spacetime curvature. Here the quantity $\zeta$ is defined as
$\zeta_{\mu\nu\alpha}=v_{\mu\alpha;\nu}-v_{\nu\alpha;\mu},$ (37)
where $v_{\mu\nu}$ appears in the Hadamard form of the Green’s function,
$G^{(1)}_{\mu\nu}(x,z)=\frac{1}{4\pi^{2}}\left(\frac{u_{\mu\nu}}{\sigma}+v_{\mu\nu}\ln|\sigma|+w_{\mu\nu}\right)\,$, (38)
and the definitions of $\sigma$, $u_{\mu\nu}$ $w_{\mu\nu}$ can be found is Refs. [Najmi1985, DeWitt1960]. In the coincidence limit one has [DeWitt1960]:
$\lim_{x\to z}v_{\mu\nu}=-\frac{1}{2}\delta^{\alpha}_{\mu}\left(R_{\alpha\nu}-\frac{1}{6}Rg_{\alpha\nu}\right).$ (39)
The non-local term (36) arises from the particle’s own retarded field and is therefore a self-interaction effect. The local Larmor contribution, in contrast, is set by the particle’s acceleration within the external MBH field. Since the test particle’s charge is presumably minuscule compared to the MBH’s magnetic charge, the contribution of the no-local term is only significant from the immediate past. We therefore evaluate Eq. (36) by its coincident limit.
For a charged particle moving on the circular orbit described in Section II, the only non-vanishing acceleration component is
$\ddot{z}^{\theta}=\pm\frac{Pq}{mr^{3}}\sqrt{\frac{\frac{M}{r}-\frac{P^{2}}{r^{2}}}{1-\frac{3M}{r}+\frac{2P^{2}}{r^{2}}}}.$ (40)
Where + stands for prograde case and - for retrograde case. Using the latter expression, one can verify that
$\lim_{x\to z}\zeta_{0\,\beta\gamma}\ddot{z}^{\gamma}=0\,$, (41)
indicating that Eq. (36) yields a negligible contribution. Analogously, one can verify that $(\dddot{z})_{0}=0$ along circular obits. Therefore, the only non-vanishing terms in Eq. (35) are
$P_{\text{tot}}=-\frac{2}{3}q^{2}\ddot{z}^{\theta}\ddot{z}_{\theta}\dot{z}_{0}-\frac{1}{3}q^{2}(R_{0\,\beta}\dot{z}^{\beta}+\dot{z}_{0}R_{\beta\gamma}\dot{z}^{\beta}\dot{z}^{\gamma})\,.$ (42)
We notice that in the limit of flat spacetime, $\dot{z}_{0}\to-1$ and the curvature contribution vanishes, recovering the familiar Larmor formula. An explicit evaluation shows that
$P_{\text{tot}}$ $=\frac{2P^{2}q^{2}\left(m^{2}+q^{2}\right)\left(\frac{M}{r}-\frac{P^{2}}{r^{2}}\right)}{3m^{2}r^{4}\sqrt{1-\frac{2M}{r}+\frac{P^{2}}{r^{2}}}}$
$\times\left(\frac{1-\frac{2M}{r}+\frac{P^{2}}{r^{2}}}{1-\frac{3M}{r}+\frac{2P^{2}}{r^{2}}}\right)^{3/2}\,.$ (43)
We are interested in assessing the stability of orbits close to the ISCO, as this is relevant for black hole environments (such as accretion disks). The energy of the particle orbiting on the ISCO is $E=-mu_{t}(r_{\text{ISCO}})\sim\mathcal{O}(1)\times m$. Thus, the characteristic radiation timescale is estimated as
$\delta\tau_{\text{rad}}\sim\frac{m}{P_{\text{tot}}(r_{\text{ISCO}})}\,.$ (44)
For values $P/M\lesssim 0.5$ and $q/m\gg 1$, Eq. (44) is approximated as
$\delta\tau_{\text{rad}}\sim 3\times 10^{4}\,{\rm s}\left(\frac{M}{10^{6}\,M_{\odot}}\right)\bigg(\frac{q}{m}\bigg)^{\!\!-4}\!\bigg(\frac{m}{M}\bigg)^{\!\!-1}\!\!\!\left(\frac{P}{M}\right)^{\!\!-2}$ (45)
For a circular orbit to be considered stable, its decay timescale must be much longer than its orbital period $t_{\text{orb}}=2\pi/\omega$, where $\omega$ is defined in Eq. (15). Evaluating the ratio $\delta\tau_{\text{rad}}/t_{\text{orb}}$ at the ISCO, we find a simple scaling law valid for values $P/M\lesssim 10^{-2}$ and $q/m\gg 1$, i.e.
$\frac{\delta\tau_{\text{rad}}}{t_{\text{orb}}}\sim 10^{2}\,\bigg(\frac{q}{m}\bigg)^{\!\!-4}\bigg(\frac{m}{M}\bigg)^{\!\!-1}\!\!\left(\frac{P}{M}\right)^{\!\!-2}$ (46)
To visualize the parameter space where orbits are both stable and non-trivial, we plot the constraints in Figure 4. The dashed lines indicates the regime where the magnetic field is strong enough to cause a significant latitudinal deviation ($|\tan\theta|<1$), as expressed in Eq. (17). The red region delineates the area of parameter space where the ratio $\delta\tau_{\text{rad}}/t_{\text{orb}}<1$, for a benchmark value of $m$ equal to the electron mass and $M=10^{6}\,M_{\odot}$. For smaller values of $m/M$, the dashed red line shifts even upward in Figure 4, whereas the opposite trend occurs if we increase $m/M$.
Figure 4: The parameter space of circular orbits for a charged particle with charge-to-mass ratio $|q/m|$ orbiting a massive black hole (MBH) with charge-to-mass ratio $|P/M|$. The dashed lines indicates where the circular orbit exhibits a significant latitudinal deviation ($|\tan\theta|<1$). The red region marks the values of the parameters for which the orbit becomes unstable under the emission of synchrotron radiation. The extent of this unstable region depends on the ratio $m/M$, as shown in Eq. (46). In this figure, we take $m$ to be the electron mass and $M=10^{6}\,M_{\odot}$.
IV Orbital structure in Rotating MBH
The charged spinning spacetime is described by the Kerr–Newman metric [Kerr1963, Mazur1982], which is given by
[…]
where
$\Delta=r^{2}-2Mr+a^{2}+P^{2},\quad\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,$ (48)
and $a=J/M$. The magnetic charge and angular momentum are constrained by
$P^{2}+a^{2}\leq M^{2}.$ (49)
As in Section II, we define the electromagnetic potential on the northern-hemisphere patch, while the southern patch is related by the gauge transformation. To derive the electromagnetic tensor in the rotating magnetically charged spacetime, we follow the complex transformation algorithm [Newman1965, Adamo2014], starting from the static spherically symmetric MBH described by the Reissner–Nordström metric.
We first replace the time coordinate $t$ with the null coordinate $u$, defined through
$u=t-\int^{r}_{0}\frac{dr^{\prime}}{1-\frac{2M}{r^{\prime}}+\frac{P^{2}}{r^{\prime 2}}},$ (50)
such that the line element in Eq. (2) simplifies to
$ds^{2}=-f(r)du^{2}-2\,du\,dr+r^{2}d\Omega^{2}.$ (51)
Then we introduce the null tetrad $\{l,n,m,\bar{m}\}$, where $l$ and $n$ are real while $m$ and $\bar{m}$ are complex conjugates. The only non-vanishing scalar products are $l\cdot n=1$ and $m\cdot\bar{m}=-1$. The contravariant metric can then be expressed in terms of the tetrad as
$g^{\mu\nu}=2(m^{(\mu}\bar{m}^{\nu)}-l^{(\mu}n^{\nu)}).$ (52)
Comparing with Eq. (51), it is easy to write down the 1-forms, and hence tetrads of the null vectors:
$\begin{split}&l=\frac{\partial}{\partial r},\\ &n=\frac{\partial}{\partial u}-\frac{1}{2}f(r)\frac{\partial}{\partial r},\\ &m=\frac{1}{\sqrt{2}r}\frac{\partial}{\partial\theta}+\frac{i}{\sqrt{2}r\sin\theta}\frac{\partial}{\partial\varphi}.\end{split}$ (53)
The 4-potential is then expressed by the null vectors as
$A_{\mu}=\frac{-iP}{\sqrt{2}r\tan\theta}(m_{\mu}-\bar{m}_{\mu}).$ (54)
To obtain the spinning solution, we complexify the Reissner–Nordström tetrad:
$\begin{split}&l\to l^{\prime}=\frac{\partial}{\partial r},\\ &n\to n^{\prime}=\frac{\partial}{\partial u}-\frac{1}{2}\left(1-\frac{M}{r}-\frac{M}{\bar{r}}+\frac{P^{2}}{|r|^{2}}\right)\frac{\partial}{\partial r},\\ &m\to m^{\prime}=\frac{1}{\sqrt{2}\bar{r}}\left(\frac{\partial}{\partial\theta}+\frac{i}{\sin\theta}\frac{\partial}{\partial\varphi}\right),\end{split}$ (55)
and simultaneously complexify the coordinates, analogously to the Schwarzschild-to-Kerr transformation:
$\begin{split}&r^{\prime}=r+ia\cos\theta,\\ &u^{\prime}=u-ia\cos\theta,\\ &\theta^{\prime}=\theta,\quad\varphi^{\prime}=\varphi.\end{split}$ (56)
It is now straightforward to express the complex tetrad in the new coordinate system (Eq. (56)) by an ordinary coordinate transformation (the primes indicating the new coordinates are dropped for simplicity):
$\begin{split}&l^{\prime}=\frac{\partial}{\partial r},\quad n^{\prime}=\frac{\partial}{\partial u}-\frac{1}{2}\left(1-\frac{2Mr-P^{2}}{\rho^{2}}\right)\frac{\partial}{\partial r},\\ &m^{\prime}=\frac{1}{\sqrt{2}(r+ia\cos\theta)}\\ &\quad\quad\left(ia\sin\theta\frac{\partial}{\partial u}-ia\sin\theta\frac{\partial}{\partial r}+\frac{\partial}{\partial\theta}+\frac{i}{\sin\theta}\frac{\partial}{\partial\varphi}\right).\end{split}$ (57)
The electromagnetic 4-potential than follows from this null tetrad via Eq. (52) to be
$A_{\mu}=\left(\frac{-Pa\cos\theta}{\rho^{2}},0,0,\frac{r^{2}+a^{2}}{\rho^{2}}P\cos\theta\right).$ (58)
which yields the covariant Maxwell tensor
$F_{\mu\nu}=\begin{bmatrix}0&-\frac{2Par\cdot\cos\theta}{\rho^{4}}&-\frac{Pa(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0\\ \frac{2Par\cdot\cos\theta}{\rho^{4}}&0&0&-\frac{2Pa^{2}r\sin^{2}\theta\cos\theta}{\rho^{4}}\\ \frac{Pa(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0&0&-\frac{P(r^{2}+a^{2})(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}\\ 0&\frac{2Pa^{2}r\sin^{2}\theta\cos\theta}{\rho^{4}}&\frac{P(r^{2}+a^{2})(r^{2}-a^{2}\cos^{2}\theta)\sin\theta}{\rho^{4}}&0\end{bmatrix}$. (59)
In a rotating charged spacetime, the behavior of particles differs between the prograde and retrograde cases. Following a similar argument as in Section II, the photon ring radius is determined by [Johannsen2013, Gal2019]
$r_{\text{pr}}^{2}-3Mr_{\text{pr}}+2P^{2}\pm 2a\sqrt{Mr_{\text{pr}}-P^{2}}=0,$ (60)
where ++ and - denote prograde and retrograde motion, respectively. For extreme MBHs, $r_{\text{ph}}$ is identical to the horizon radius $r_{\text{h}}=M$.
The ISCO radius is obtained by defining [Liu2017]
$\begin{split}\mathcal{R}(r):=&\rho^{4}(\frac{dr}{d\tau})^{2}\\ =&\pm 2aj\varepsilon(P^{2}-2Mr)-(\Delta-a^{2})(r^{2}+j^{2})\\ &+r^{4}\varepsilon^{2}+a^{2}\left(\varepsilon^{2}(r^{2}+2Mr-P^{2})-r^{2}\right),\end{split}$ (61)
where ++ and - denote prograde and retrograde motion, $\varepsilon$ and $j$ denote the energy and angular momentum per unit mass. Stable circular orbits satisfy
$\mathcal{R}(r)=\frac{d\mathcal{R}(r)}{dr}=0,\quad\frac{d^{2}\mathcal{R}(r)}{dr^{2}}\geq 0,$ (62)
which yields the ISCO condition
$\begin{split}Mr(6Mr-r^{2}-9P^{2}+3a^{2})&+4P^{2}(P^{2}-a^{2})\\ \mp 8a(Mr-P^{2})^{3/2}&=0.\end{split}$ (63)
Here - and ++ denote prograde and retrograde motion, respectively.
The electromagnetic field in this spacetime diverts electrically charged particles from the equator in a similar way. While the electromagnetic field in this spacetime is complicated, we again have off-equatorial circular orbits similar to those discussed in Section II.
We begin by examining the $t$ and $\varphi$ components of the geodesic equations Eq. (6). It turns out that the relationships in Eq. (13) still hold for a rotating MBH. Consequently, for circular orbits with $u^{r}=u^{\theta}=0$, all terms in these equations vanish.
Figure 5: $\theta(r)$ of a q/m=1q/m=-1 particle prograding a $P/M=0.6$ spinning MBH, for multiple $P/M$ values. The peak and smallest allowed radius decrease monotonously with $a/M$.
The remaining two equations determine the angular velocity and polar angle of a particle on a spherical orbit. Due to their complexity, the two variables, $\theta$ and $\omega$, are coupled in the equations of motion. Their explicit forms are relegated to Appendix A and can only be solved numerically. It is worth noting that the Carter constant can be used to decouple the radial and polar equations [Dyson2023]. Nevertheless, this approach introduces offsetting complexities into the temporal and azimuthal components of the motion. We therefore proceed with the direct numerical solution of the coupled system.
Because of the periodicity of the trigonometric functions involved, these equations yield infinitely many solutions. However, within the range $0\leq\theta\leq\pi$, only two distinct branches exist, corresponding to prograde and retrograde particles. These two branches are shown in Figure 5 (prograde) and Figure 6 (retrograde), respectively. In each plot, we fix the particle charge-to-mass ratio at q/m=1q/m=-1, set the magnetic charge to $P=0.6M$, and vary the rotation parameter from $a=0$ to $a=0.8M$, approaching the extremal case.
Figure 6: $\theta(r)$ of a q/m=1q/m=-1 particle retrograding a $P/M=0.6$ spinning MBH, for multiple $P/M$ values. The peak and smallest allowed radius increase monotonously with $a/M$.
The behavior of $\theta(r)$ is qualitatively similar to the non-rotating case: it approaches $\pi/2$ in both the small- and large-radius limits, with a single intermediate peak. An exception arises for prograde motion around an extremal MBH, where $\theta$ attains its extremum near the smallest admissible radius. For retrograde orbits, both the peak and the innermost allowed radius occur farther from the black hole, consistent with the intuitive picture that prograde motion is more closely aligned with the spacetime geometry. Contrary to Section II, the ISCO is not well defined for charged particles in this spacetime, since $\mathcal{R}(r)$ depends on three constants, energy, angular momentum, and the Carter constant [Dyson2023], thus the ISCO condition (Eq. (62)) does not yield a unique marginal orbit. Consequently, the peak radius lacks a direct physical interpretation.
Despite this, the photon ring still serves as the absolute inner boundary for circular orbits. To demonstrate it, we analyze the radial component of the equation of motion, which determines the radial acceleration $R(r,\theta,\omega)=d^{2}r/d\tau^{2}$. A circular orbit can only exist if there is a set of parameters $(\theta,\omega)$ for which $R=0$. We numerically computed the maximum possible value of $R$ over all latitudes and a wide range of angular velocities at a given radius $r$.
Figure 7: Maximum radial acceleration $R(r,\theta,\omega)$ over all latitudes and $M\omega\leq 10$, as a function of the relative radius to photon ring radius $r/r_{\text{pr}}$. The black dashed lines mark $r=r_{\text{pr}}$ and $R=0$. A sharp transition occurs at the photon ring radius $r=r_{\text{pr}}$: $\max(R)<0$ for $r<r_{\text{pr}}$, prohibiting circular motion, while $\max(R)>0$ for $r>r_{\text{pr}}$, allowing circular orbits.
The results are shown in Figure 7. There is a sharp transition at the photon ring radius ($r=r_{\text{pr}}$). For all radii inside the photon ring ($r<r_{\text{pr}}$), the maximum value of $R$ is negative. This implies that no amount of orbital velocity can counteract the gravitational pull, and any particle is destined to plunge. For $r>r_{\text{pr}}$, $\max(R)$ becomes positive, allowing for solutions where $R=0$. Thus, the photon ring marks the critical inner boundary for the existence of off-equatorial circular orbits for charged particles in the rotating spacetime.
In Figure 7 we show the maximum of $R(\theta,\omega)$ in the range of $0\leq\theta\leq\pi,\quad 0\leq M\omega\leq 10$ as a function of the relative radius to photon ring radius ($r/r_{\text{pr}}$), for prograde and retrograde test particle with q/m=1q/m=-1, for $a/M$ from 0 to extremum (denoted by color from purple to yellow) at three different $P/M$. There is a jump point at $r_{\text{pr}}$: inside the photon ring radius, the maximum of $R(\theta,\omega)$ is negative, thus excluding any solution of a circular orbit; when $r>r_{\text{pr}}$ however, the maximum depends on the range of $\omega$ but tends to infinity as $\omega$ increases.
We then conclude that $R(\theta,\omega)$ has no root at $r<r_{\text{pr}}$, and from Section IV, we see that Eq. (6) always have solution. Thus, the photon ring radius marks the critical radius above which there exist off-equatorial circular orbits for charged particles.
Finally, it is worth noting that rotating electrically charged black holes also possess radial and polar components of both the electric and magnetic field. However, in contrast to the magnetically charged case, no off-equatorial circular orbits exist here. For an electrically charged particle, circular motion is confined to the equatorial plane only.
Figure 8: $\Theta(\theta,\omega,r)$ in Kerr–Newman space-time with $a/M=Q/M=0.5$, evaluated at a wide range of $\omega$, and three radii: the photon ring ($r=r_{\text{pr}}$), the ISCO, and a large-$r$ regime, for a prograde q/m=1q/m=-1 particle. At all radii, $\Theta$ vanishes only in the equatorial limit $\theta\to\pi/2$. A secondary trend is visible: near the photon ring, $\Theta$ is suppressed at large $\omega$, while at large radii it is suppressed at small $\omega$, reflecting the shift in angular motion with radius.
The electromagnetic tensor in Kerr–Newman spacetime takes the form
$F^{(KN)}_{\mu\nu}=\begin{bmatrix}0&-\frac{Q(r^{2}-a^{2}\cos^{2}\theta)}{\rho^{4}}&\frac{2Qra^{2}\sin\theta\cos\theta}{\rho^{4}}&0\\ \frac{Q(r^{2}-a^{2}\cos^{2}\theta)}{\rho^{4}}&0&0&-\frac{Qa\sin^{2}\theta(r^{2}-a^{2}\cos^{2}\theta)}{\rho^{4}}\\ -\frac{2Qra^{2}\sin\theta\cos\theta}{\rho^{4}}&0&0&\frac{2Qra(r^{2}+a^{2})\sin\theta\cos\theta}{\rho^{4}}\\ 0&\frac{Qa\sin^{2}\theta(r^{2}-a^{2}\cos^{2}\theta)}{\rho^{4}}&\frac{2Qra(r^{2}+a^{2})\sin\theta\cos\theta}{\rho^{4}}&0\end{bmatrix}$. (64)
Both $F^{(KN)}_{t\theta}$ and $F^{(KN)}_{\theta\varphi}$ vanish at $\theta=\pi/2$, since they contain an overall factor of $\cos\theta$. Thus, on the equatorial plane, only a radial electric field and a polar magnetic field remain, yielding a purely radial electromagnetic force that admits equatorial orbits. When $\theta\neq\pi/2$, however, the $\theta$-component of Eq. (6) has no physical solution for $\theta$ and $\omega$, thereby excluding off-equatorial circular orbits.
To show this explicitly, we define
$\Theta(\theta,\omega,r)=d^{2}\theta/d\tau^{2}\,$, (65)
as the total polar acceleration. Figure 8 plots $\Theta$ for a charged particle in an electrically charged KN spacetime. Away from the equatorial plane, $\Theta$ is strictly non-zero, confirming that no equilibrium solution exists. This highlights a fundamental difference: the purely radial field of a magnetic monopole provides a unique mechanism for latitudinal confinement that is absent in the field of a rotating electric charge.
V Conclusions
In conclusion, we have presented a systematic and detailed analysis of the orbital dynamics of charged test particles around both static and rotating magnetically charged black holes. Although the motion of neutral particles in such spacetimes is well understood, we have demonstrated that the interplay between the gravitational field and the long-range magnetic Lorentz force gives rise to a novel family of stable, off-equatorial circular orbits. These trajectories, confined to a fixed latitude, represent a significant departure from the planar motion typically associated with black hole orbits.
Our primary results were derived for the static, spherically symmetric MBH. We obtained a closed-form analytical expression, Eq. (16), which determines the equilibrium polar angle $\theta$ of a charged particle as a function of its orbital radius. Through a rigorous analysis, we established profound connections between these new orbits and the fundamental structure of the spacetime including the photon sphere and ISCO. Our analysis of synchrotron radiation confirms these orbits are viable over astrophysical timescales. We found that even electrons, which maximizes radiative losses, can maintain a stable orbit with a significant latitudinal deviation.
We extended this investigation to rotating MBH spacetime, where we numerically computed the prograde and retrograde orbital branches. Our results show that the MBH spin significantly modifies the orbits, pulling prograde trajectories closer and pushing retrograde ones farther away. Despite the added complexity and the loss of a direct ISCO connection, we demonstrated that the photon ring robustly maintains its role as the inner boundary for stable circular motion. We also highlighted a crucial distinction by showing that such off-equatorial orbits are forbidden in the spacetime of an electrically charged Kerr-Newman black hole, underscoring the unique role of the radial magnetic field in providing latitudinal confinement.
The existence of these stable, nonplanar orbits opens several avenues for future work. They may produce unique observational signatures, such as warps or gaps in accretion disks, or distinctive polarization patterns in the emitted synchrotron radiation. Future theoretical studies could explore the stability of these orbits against radial and vertical perturbations, include the effects of particle spin, or perform a more detailed analysis of the radiation reaction force. This work provides a clear and robust foundation for understanding the rich dynamics of charged matter in the vicinity of magnetically charged compact objects.
VI Acknowledgments
The authors gratefully acknowledge Julian Krolik and Emanuele Berti for their stimulating discussions. We thank David Pereñiguez for his insightful comments on the manuscript. Xilai Li would like to thank the Shanghai Jiao Tong University International Cooperation and Exchange Office and the Johns Hopkins University Global Education Office for facilitating a visit to JHU. L.D.G. is supported by NSF Grants No. PHY-2207502, AST-2307146, PHY-090003 and PHY-20043, by NASA Grant No. 21-ATP21-0010, by the John Templeton Foundation Grant 62840, and by the Simons Investigator Grant No. 144924.
Appendix A Equations of Motion of Charged Particles Around a Rotating MBH
Consider a charged test particle moving on a circular orbit with 4-velocity
$u^{\mu}=(u^{t},0,0,\omega).$
The normalization condition $u^{\mu}u_{\mu}=-1$ fixes $u^{t}$ as a function of $\omega$:
$u^{t}=\frac{-g_{t\varphi}\omega+\sqrt{(g_{t\varphi}\omega)^{2}-g_{\varphi\varphi}\omega^{2}-1}}{g_{tt}},$ (66)
where the metric coefficients are given in Eq. (47).
Expanding the radial component of the equation of motion with both the Christoffel symbols and the electromagnetic tensor Eq. (59), one finds the $r$-equation
$\begin{split}\frac{d^{2}r}{d\tau^{2}}=&\frac{\left(P^{2}+r(r-2M)+1\right)\left(\cos(2\theta)M+M+2r\left(P^{2}-Mr\right)\right)(u^{t})^{2}}{2\left(r^{2}+\cos^{2}\theta\right)^{3}}\\ -&\frac{\left(P^{2}+r(r-2M)+1\right)\omega\left(M\cos^{2}\theta+r\left(P^{2}-Mr\right)\right)u^{t}\sin^{2}\theta}{\left(r^{2}+\cos^{2}\theta\right)^{3}}\\ -&\frac{4\left(P^{2}+r(r-2M)+16\right)\omega\left(16M\cos^{2}\theta+r\left(P^{2}-Mr\right)\right)u^{t}\sin^{2}\theta}{\left(r^{2}+16\cos^{2}\theta\right)^{3}}\\ +&\frac{\left(P^{2}+r(r-2M)+16\right)\omega^{2}\left(\left(256(M-r)\sin^{2}\theta+r\left(r^{4}+16\left(P^{2}-Mr+16\right)\sin^{2}\theta-256\right)+32r\left(r^{2}+16\right)\right)\cos^{2}\theta\right)}{\left(r^{2}+16\cos^{2}\theta\right)^{3}\csc^{2}\theta}\\ +&\frac{q}{m\rho^{2}}\Biggl[a\left(2Mr-P^{2}\right)u^{t}\Biggl(\frac{2a^{2}Pr\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta\left(\cos(2\theta)a^{2}+a^{2}+2\left(P^{2}+r^{2}-2Mr\right)\right)}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\\ &\quad\quad\quad-\frac{4a^{2}Pr\left(P^{2}-2Mr\right)\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\Biggr)\sin^{2}\theta\\ &\quad+a\left(2Mr-P^{2}\right)\omega\Biggl(\frac{4Pr\left(P^{2}-2Mr\right)\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta\sin^{2}\theta a^{3}}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\\ &\quad\quad\quad+\frac{8Pr\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta\left(\left(a^{2}+r^{2}\right)^{2}-a^{2}\left(a^{2}+P^{2}+r^{2}-2Mr\right)\sin^{2}\theta\right)a}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{2}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)^{2}}\Biggl)\sin^{2}\theta\\ &\quad-\omega\Biggl(\frac{2a^{2}Pr\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta\left(\cos(2\theta)a^{2}+a^{2}+2\left(P^{2}+r^{2}-2Mr\right)\right)}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\\ &\quad\quad\quad-\frac{4a^{2}Pr\left(P^{2}-2Mr\right)\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\Biggl)\\ &\quad\quad\quad\quad\times\left(\left(a^{2}+r^{2}\right)^{2}\sin^{2}\theta-a^{2}\left(a^{2}+P^{2}+r^{2}-2Mr\right)\sin^{4}\theta\right)\\ &\quad+u^{t}\left(a^{2}\cos^{2}\theta+P^{2}+r^{2}-2Mr\right)\Biggl(\frac{4Pr\left(P^{2}-2Mr\right)\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta\sin^{2}\theta a^{3}}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\\ &\quad\quad\quad+\frac{8Pr\left(a^{2}+P^{2}+r^{2}-2Mr\right)\cos\theta\left(\left(a^{2}+r^{2}\right)^{2}-a^{2}\left(a^{2}+P^{2}+r^{2}-2Mr\right)\sin^{2}\theta\right)a}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{2}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)^{2}}\Biggl)\Biggl]=0,\end{split}$ (67)
and the $\theta$-equation
$\begin{split}\frac{d^{2}\theta}{d\tau^{2}}=&\frac{\left(P^{2}-2Mr\right)\left(r^{2}+1\right)\omega\cos\theta\sin\theta u^{t}}{\left(r^{2}+\cos^{2}\theta\right)^{3}}-\frac{\left(P^{2}-2Mr\right)\cos\theta\sin\theta(u^{t})^{2}}{\left(r^{2}+\cos^{2}\theta\right)^{3}}+\frac{4\left(P^{2}-2Mr\right)\left(r^{2}+16\right)\omega\cos\theta\sin\theta u^{t}}{\left(r^{2}+16\cos^{2}\theta\right)^{3}}\\ -&\frac{16\omega^{2}\cos\theta\sin^{3}\theta\left(16\left(P^{2}+r^{2}-2Mr+16\right)\sin^{2}\theta-\left(r^{2}+16\right)^{2}\right)}{\left(r^{2}+16\cos^{2}\theta\right)^{3}}\\ +&\frac{\omega^{2}\left(r^{2}+16\cos^{2}\theta\right)\left(\left(r^{2}+16\right)^{2}-32\left(P^{2}+r^{2}-2Mr+16\right)\sin^{2}\theta\right)\sin(2\theta)}{2\left(r^{2}+16\cos^{2}\theta\right)^{3}}\\ +&\frac{q}{m\rho^{2}}\Biggl[a\left(2Mr-P^{2}\right)u^{t}\Biggl(\frac{P\left(a^{2}+r^{2}\right)\left(r^{2}-a^{2}\cos^{2}\theta\right)\left(\cos(2\theta)a^{2}+a^{2}+2\left(P^{2}+r^{2}-2Mr\right)\right)\csc\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\\ &\quad\quad\quad-\frac{2a^{2}P\left(P^{2}-2Mr\right)\left(r^{2}-a^{2}\cos^{2}\theta\right)\sin\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\Biggl)\sin^{2}\theta\\ &\quad+a\left(2Mr-P^{2}\right)\omega\Biggl(\frac{4aP\left(r^{2}-a^{2}\cos^{2}\theta\right)\left(\left(a^{2}+r^{2}\right)^{2}-a^{2}\left(a^{2}+P^{2}+r^{2}-2Mr\right)\sin^{2}\theta\right)\sin\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{2}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)^{2}}\\ &\quad\quad\quad+\frac{2aP\left(P^{2}-2Mr\right)\left(a^{2}+r^{2}\right)\left(r^{2}-a^{2}\cos^{2}\theta\right)\sin\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\Biggl)\sin^{2}\theta\\ &\quad-\omega\left(\left(a^{2}+r^{2}\right)^{2}\sin^{2}\theta-a^{2}\left(a^{2}+P^{2}+r^{2}-2Mr\right)\sin^{4}\theta\right)\\ &\quad\quad\times\Biggl(\frac{P\left(a^{2}+r^{2}\right)\left(r^{2}-a^{2}\cos^{2}\theta\right)\left(\cos(2\theta)a^{2}+a^{2}+2\left(P^{2}+r^{2}-2Mr\right)\right)\csc\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\\ &\quad\quad\quad-\frac{2a^{2}P\left(P^{2}-2Mr\right)\left(r^{2}-a^{2}\cos^{2}\theta\right)\sin\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\Biggl)\\ &\quad+u^{t}\left(a^{2}\cos^{2}\theta+P^{2}+r^{2}-2Mr\right)\Biggl(\frac{4aP\left(r^{2}-a^{2}\cos^{2}\theta\right)\left(\left(a^{2}+r^{2}\right)^{2}-a^{2}\left(a^{2}+P^{2}+r^{2}-2Mr\right)\sin^{2}\theta\right)\sin\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{2}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)^{2}}\\ &\quad\quad\quad+\frac{2aP\left(P^{2}-2Mr\right)\left(a^{2}+r^{2}\right)\left(r^{2}-a^{2}\cos^{2}\theta\right)\sin\theta}{\left(a^{2}+P^{2}+r(r-2M)\right)\left(r^{2}+a^{2}\cos^{2}\theta\right)^{3}\left(\cos(2\theta)a^{2}+a^{2}+2r^{2}\right)}\Biggl)\Biggl]=0.\end{split}$ (68)
Where $\rho$ is defined in Eq. (48) and the explicit form of $u^{t}$ is presented above in Eq. (66). Together, the $r$- and $\theta$-equations determine admissible pairs $(\theta,\omega)$ at each orbital radius. For every radius larger than the photon-ring radius, the equations yield two physical solutions for $\omega$: one positive (prograde motion) and one negative (retrograde motion).
In Kerr-Newman spacetime the equations take a similar form. However, all circular solutions are confined to $\theta=\pi/2$.